a three resistors of 4ohms, 12 ohms and 5 ohms are connected in parallel. if total current is 12A, find the current through each resistor


1
Expert's answer
2021-12-15T11:32:06-0500

Total resistance "R_T=\frac{1}{R_1}+\frac{1}{R_2}+\frac{1}{R_3}"

"\frac{1}{R_T}=\frac{1}{4}+\frac{1}{12}+\frac{1}{5}= \frac{8}{15}"

"R_T=1.87ohms"

V=IR

V= 12×1.87=22.5V

"I_1=\frac{22.5}{4}=5.625A"

"I_2=\frac{22.5}{12}=1.875A"

"I_3=\frac{22.5}{5}=4.5A"